OpenStudy (anonymous):

A local snail population grows according to the function g(x) = 200(1.03)2x. Demonstrate the steps to convert g(x) into an equivalent function with only x as the exponent. Then, explain to Iris how the key features of this local snail population compares to the key features of the invasive population.

OpenStudy (anonymous):

@ninjasandtigers

OpenStudy (anonymous):

\[g(x)=200(1.03)^{2x}\] \[g(x)=200((1.03)^2)^x\] \[g(x)=200(1.0609)^x\]

OpenStudy (anonymous):

Can you help me with 2 more on 2 different threads?

OpenStudy (anonymous):

I would love to but I have to take my calculus test that I also need help on

OpenStudy (anonymous):

Okay, thank you so much for help.

OpenStudy (anonymous):

welcome :)